The probability of getting 11 with one throw of 2 dice is 1/6*1/6*2 = 1/18
So the probability of not getting 11 with 1 throw of the dice is 17/18.
Tossing the dice 54 times, the probability of not getting 11 54 times is (17/18)54 = 0.0456...
So the probability of at least 1 roll of 11 is 1 - 0.0456 = 0.954
Are any US coins more likely to land on heads or tails?
No there is a fifty fifty chance of getting heads or tails
There are 11 crayons, so the chance of picking any specific crayon is 1/11. Because four of them are red, the chance of choosing one of the four red crayons is 4/11.
After the first draw, there are 10 crayons remaining, three of which are red (assuming the first crayon drawn was red; if it was some other color, we don't care, because there's then no chance that we will end up with two red crayons). So the chance of drawing a red crayon on the second draw is 3/10.
The overall chance of pulling two red crayons in a row is 4/11 x 3/10, or 12/110 ... a shade under 11 percent.
Consider the word Mississippi find the total number of distinct permutations of its letters?
There are 11 letters in all, so there are a total of 11! = 39,916,800 permutations. The letter 'i' appears 4 times, so every distinct permutation will be duplicated 4! = 24 times, so we divide the total by 24. Similarly, the letter 's' appears 4 times so we divide again by 24. Finally, we divide by two because half the remaining permutations have transposed the letter p. Thus we get 11! / (4! * 4! * 2!) = 34,650 distinct permutations.
No.
The process is to add or subtract 0.5, (not 5), from the boundaries of the interval. Strictly speaking, the amount to be added (or subtracted) is half the measurement unit. Many discrete variables are measured in integers but:
UK shoe sizes are measured in half-numbers. So you have sizes 7, 71/2, 8, 81/2 and so on. The continuity correction would be 0.25.
What is the probability of drawing a diamond from a normal deck of cards?
If you draw 40 cards without replacement the probability is 1!
If you draw just one, the probability is 1/4.
Mia has 20 marbles Sixteen of them are red What fraction of the marbles are red?
16/20, which can be simplified if you so wish.
Why Central Limit Theorem rolling a dice once you will never have a bell shaped curve?
Rolling a die once will give one point. A single point does not even define a line, let alone a curve!
What are the 21 combinations when you roll two dice?
11, 12, 13, 14, 15, 16
22, 23, 24, 25, 26,
33, 34, 35, 36,
44, 45, 46,
55, 56,
66.
6*365/2 times = 1095 times
no
What is importance of central limit theorem?
The importance is that the sum of a large number of independent random variables is always approximately normally distributed as long as each random variable has the same distribution and that distribution has a finite mean and variance. The point is that it DOES NOT matter what the particular distribution is. So whatever distribution you start with, you always end up with normal.
1/6 * 1/6 = 1/36
What is the probability of a head on the first flip and a tail on the second flip?
Probability of H on the first flip = 1/2
Probability of T on the second flip = 1/2
Probability of both = (1/2 x 1/2) = 1/4 = 25%
if we remove those 4 cards then we have the cards from 2,3,4,5,6,7,8,9,10
use arithmetic progression with first and last elements i.e a,l
sum of all digits =n/2(a+l);
n=9
a=2
l=10
sum=9/2(2+10)
=9/2(12)
=54
but 4 times = 4*54=216
Why normal distribution is called normal distribution?
Most random variables are found to follow the probability distribution function
All this means is that most things which can be measured quantitatively, like a population's height, the accuracy of a machine, effectiveness of a drug on fighting bacteria, etc. will occur with a probability that can be calculated according to this equation. Since most things follow this equation, this equation is considered to be the "normal" probability density. "Normal" events follow a "normal" probability distribution.
9 X 8 = 72: For the winner's first choice there are nine alternatives, and for the second choice there are only eight alternatives.
A fair die (fair dice) is a dice which isn't altered in any way, this means it has the same chance of landing on any of one of the faces.
The opposite to this is a loaded dice or cheat dice. These dice have been altered in some way, maybe by changing the weight distribution so that one of the numbers on the die/dice are more likely to be rolled.
It is 60/100 = 0.6
How many outcomes can you get if a code contains 3 numbers and you can use 1-9?
Since the code contains three digits we can set up a template like this: _ _ _. Since there are 9 numbers between 1 and 9 inclusive we just do 9*9*9 = 729.